# Parabolic induction in the homotopy category of pro-$p$ Iwahori-Hecke modules

### v2

@@ 5.2. Supersingular characters in ${\operatorname{Ho}}({\mathcal{H}_{\operatorname{aff}}})$
Similarly to the previous example we have decompositions ${\operatorname{Res}}^{\mathcal{H}}_{{\mathcal{H}}^{\dagger}_{x}}({\mathfrak{m}})\cong{\mathfrak{m}}_{0}\oplus{\mathfrak{m}}_{1}\oplus{\mathfrak{m}}_{2}$ and ${\operatorname{Res}}^{\mathcal{H}}_{\mathcal{H}_{\operatorname{aff}}}({\mathfrak{m}}^{\prime})\cong{\mathfrak{m}}^{\prime}_{0}\oplus{\mathfrak{m}}^{\prime}_{1}\oplus{\mathfrak{m}}^{\prime}_{2}$. Here, all ${\mathfrak{m}}_{j}$ and ${\mathfrak{m}}^{\prime}_{j}$ are one-dimensional, and ${\mathcal{H}}_{x}$ acts on ${\mathfrak{m}}_{0}$, ${\mathfrak{m}}_{1}$ and ${\mathfrak{m}}_{2}$ via the characters $(-1,0,-1)_{\xi}$, $(-1,-1,0)_{\xi}$ and $(0,-1,-1)_{\xi}$ respectively, and on ${\mathfrak{m}}^{\prime}_{0}$, ${\mathfrak{m}}^{\prime}_{1}$ and ${\mathfrak{m}}^{\prime}_{2}$ via the characters $(0,0,-1)_{\xi}$, $(-1,0,0)_{\xi}$ and $(0,-1,0)_{\xi}$ respectively. Furthermore, $\hat{\omega}^{3}$ acts as $\lambda{\operatorname{id}}$ on both ${\mathfrak{m}}$ and ${\mathfrak{m}}^{\prime}$. Thus we see that ${\mathfrak{m}}_{0}\cong{\mathfrak{m}}^{\prime}_{0}$ and ${\mathfrak{m}}_{1}\cong{\mathfrak{m}}^{\prime}_{2}$, and 5.10(ii) together with [19, Lemma 3.4] imply that ${\mathfrak{m}}_{2}$ and ${\mathfrak{m}}^{\prime}_{1}$ have finite projective dimension over ${\mathcal{H}}^{\dagger}_{x}$. Putting everything together, this gives
@@ 5.3. The $\widetilde{\Omega}$-action on morphisms
We recall the action of subgroups of $\Omega$ on ${\operatorname{Hom}}$ spaces for various subalgebras of ${\mathcal{H}}$ as in [3, §4.2]. Let $A,B\subseteq{\mathcal{H}}$ be two Gorenstein $k$-subalgebra such that $k[T(\mathbb{F}_{q})]\subseteq A\subseteq B$ and $B$ is free over $A$, both as a left and as a right module. Furthermore, we also let $\widetilde{\Omega}_{B}:=\widetilde{\Omega}\cap B$ and denote by $\widetilde{\Omega}_{A}$ the normaliser of $A$ in $\widetilde{\Omega}_{B}$. One example to have in mind is $A={\mathcal{H}_{\operatorname{aff}}}$, $\widetilde{\Omega}_{A}=\widetilde{\Omega}_{B}=\widetilde{\Omega}$ and $B={\mathcal{H}}$, but we will see other situations where this applies as well. We will write by $\Omega_{A}$ the image of $\widetilde{\Omega}_{A}$ under the canonical surjection $\widetilde{\Omega}\twoheadrightarrow\Omega$. We also denote by ${\mathcal{H}}_{A,\Omega_{A}}$ the $k$-subalgebra of ${\mathcal{H}}$ which is generated by $A$ and $\{T_{\tilde{\omega}}\}_{\tilde{\omega}\in\widetilde{\Omega}_{A}}$. To simplify notation, for any ${\mathfrak{m}}\in{\operatorname{Mod}}(B)$ we will denote ${\operatorname{Res}}^{B}_{A}({\mathfrak{m}})$ by ${\mathfrak{m}}|_{A}$.
@@
We fix $\tilde{\omega}\in\widetilde{\Omega}_{B}$ and let $A^{\tilde{\omega}}:=T_{\tilde{\omega}^{-1}}AT_{\tilde{\omega}}\subseteq B$. We have $A^{\tilde{\omega}}=A$ whenever $\tilde{\omega}\in\widetilde{\Omega}_{A}$, and in general the map $a\mapsto T_{\tilde{\omega}^{-1}}aT_{\tilde{\omega}}$ defines an isomorphism $A\cong A^{\tilde{\omega}}$ of $k$-algebras. Thus $A^{\tilde{\omega}}$ is also Gorenstein. This isomorphism induces an exact equivalence of categories ${\operatorname{Mod}}(A)\simeq{\operatorname{Mod}}(A^{\tilde{\omega}})$, $\mathfrak{n}\mapsto\mathfrak{n}^{\tilde{\omega}}$, where $\mathfrak{n}^{\tilde{\omega}}$ is defined to have the same underlying $k$-vector space as $\mathfrak{n}$ with $A^{\tilde{\omega}}$-action given by $x\cdot(T_{\tilde{\omega}^{-1}}aT_{\tilde{\omega}}):=xa$ for all $a\in A$ and all $x\in\mathfrak{n}^{\tilde{\omega}}$. Note that this equivalence preserves free and projective modules, and hence preserves Gorenstein projective modules as well by exactness. In addition, we observe that if ${\mathfrak{m}}\in{\operatorname{Mod}}(B)$ then the map $x\mapsto xT_{\tilde{\omega}}$ defines a natural $A^{\tilde{\omega}}$-linear isomorphism $({\mathfrak{m}}|_{A})^{\tilde{\omega}}\cong{\mathfrak{m}}|_{A^{\tilde{\omega}}}$. By putting ${\mathfrak{m}}=B$, this implies that $B$ is also free over $A^{\tilde{\omega}}$.
@@
Suppose now that ${\mathfrak{m}}_{1},{\mathfrak{m}}_{2}\in{\operatorname{Mod}}(B)$ and let $f\in{\operatorname{Hom}}_{A}({\mathfrak{m}}_{1}|_{A},{\mathfrak{m}}_{2}|_{A})$. By considering the image of $f$ under the above equivalence of module categories and pulling back along the isomorphisms $({\mathfrak{m}}_{i}|_{A})^{\tilde{\omega}}\cong{\mathfrak{m}}_{i}|_{A^{\tilde{\omega}}}$ ($i=1,2$), we obtain an $A^{\tilde{\omega}}$-linear map $f\cdot\tilde{\omega}:{\mathfrak{m}}_{1}|_{A^{\tilde{\omega}}}\to{\mathfrak{m}}_{2}|_{A^{\tilde{\omega}}}$ defined by $(f\cdot\tilde{\omega})(x)=f(xT_{\tilde{\omega}^{-1}})T_{\tilde{\omega}}$ for $x\in{\mathfrak{m}}_{1}$. Thus we have constructed a map
@@
The above construction extends to morphisms in the homotopy category as follows. Assume that ${\mathfrak{m}}_{1}|_{A}$ is Gorenstein projective and let $\theta:\mathfrak{p}\twoheadrightarrow{\mathfrak{m}}_{2}$ be a surjection from a projective $B$-module. Note that ${\mathfrak{m}}_{1}|_{A^{\tilde{\omega}}}\cong({\mathfrak{m}}_{1}|_{A})^{\tilde{\omega}}$ is also Gorenstein projective, that ${\operatorname{Res}}^{B}_{A}(\theta)\cdot\tilde{\omega}={\operatorname{Res}}^{B}_{A^{\tilde{\omega}}}(\theta)$, and that $\mathfrak{p}|_{A}$ and $\mathfrak{p}|_{A^{\tilde{\omega}}}$ are also projective, where the latter holds because $B$ is free over $A$ and $A^{\tilde{\omega}}$. Moreover, the fact that $(-)\cdot\tilde{\omega}$ is compatible with composition implies that the square
[…]
commutes. Since the images of the two vertical arrows above are precisely those maps which factor through a projective module (c.f. 2.15), it then follows from 2.14 that $(-)\cdot\tilde{\omega}$ descends to a map $[{\mathfrak{m}}_{1},{\mathfrak{m}}_{2}]_{A}\to[{\mathfrak{m}}_{1},{\mathfrak{m}}_{2}]_{A^{\tilde{\omega}}}$. In particular, the $\widetilde{\Omega}_{A}$-action from above descends to an action on $[{\mathfrak{m}}_{1},{\mathfrak{m}}_{2}]_{A}$.
Next, assume that ${\mathfrak{m}}_{1}$ is Gorenstein projective in ${\operatorname{Mod}}(B)$. Then by 2.14 and the above, we have a commutative diagram
[…]
where the top row is exact. From this we deduce the existence of the dashed map $[{\mathfrak{m}}_{1},{\mathfrak{m}}_{2}]_{B}\to[{\mathfrak{m}}_{1},{\mathfrak{m}}_{2}]_{A}^{\Omega_{A}}$. After composing with the inclusion $[{\mathfrak{m}}_{1},{\mathfrak{m}}_{2}]_{A}^{\Omega_{A}}\subseteq[{\mathfrak{m}}_{1},{\mathfrak{m}}_{2}]_{A}$, this is by construction just the map induced by the functor ${\operatorname{Ho}}({\operatorname{Res}}^{B}_{A})$. This generalises as follows:
@@
Suppose that ${\mathfrak{m}}_{1},{\mathfrak{m}}_{2}\in{\operatorname{Mod}}(B)$ and that ${\mathfrak{m}}_{1}|_{A}$ is Gorenstein projective, and let $\tilde{\omega}\in\widetilde{\Omega}_{B}$. Then the diagram
[…]
@@
Proof.
Given a $B$-linear map $f:{\mathfrak{m}}_{1}\to{\mathfrak{m}}_{2}$, we write $\overline{f}\in[{\mathfrak{m}}_{1},{\mathfrak{m}}_{2}]_{B}$ for the image of $f$ in ${\operatorname{Ho}}(B)$. Assuming first that ${\mathfrak{m}}_{1}$ is Gorenstein projective, we then have that any $g\in[{\mathfrak{m}}_{1},{\mathfrak{m}}_{2}]_{B}$ is equal to $\overline{f}$ for some $f$. The commutativity of (5.10) in that case follows from the fact that ${\operatorname{Res}}^{B}_{A}(f)\cdot\tilde{\omega}={\operatorname{Res}}^{B}_{A^{\tilde{\omega}}}(f)$, which holds because $f$ is $B$-linear and hence $\widetilde{\Omega}_{B}$-equivariant.
In general, let $q_{c}:Q_{c}{\mathfrak{m}}_{1}\to{\mathfrak{m}}_{1}$ be a Gorenstein projective replacement in ${\operatorname{Mod}}(B)$ and pick $g\in[{\mathfrak{m}}_{1},{\mathfrak{m}}_{2}]_{B}$. Note that ${\operatorname{Res}}^{B}_{A}(q_{c})\cdot\tilde{\omega}={\operatorname{Res}}^{B}_{A^{\tilde{\omega}}}(q_{c})$ is also a Gorenstein projective replacement in ${\operatorname{Mod}}(A^{\tilde{\omega}})$. From the Gorenstein projective case above, the functoriality of ${\operatorname{Ho}}({\operatorname{Res}}^{B}_{A})$ and ${\operatorname{Ho}}({\operatorname{Res}}^{B}_{A^{\tilde{\omega}}})$, and the fact that $(-)\cdot\tilde{\omega}$ is compatible with composition, we now get
$({\operatorname{Ho}}({\operatorname{Res}}^{B}_{A})(g)\cdot\tilde{\omega})\circ{\operatorname{Ho}}({\operatorname{Res}}^{B}_{A^{\tilde{\omega}}})(q_{c})$ $=({\operatorname{Ho}}({\operatorname{Res}}^{B}_{A})(g)\cdot\tilde{\omega})\circ({\operatorname{Ho}}({\operatorname{Res}}^{B}_{A})(q_{c})\cdot\tilde{\omega})$
$={\operatorname{Ho}}({\operatorname{Res}}^{B}_{A})(g\circ q_{c})\cdot\tilde{\omega}$
$={\operatorname{Ho}}({\operatorname{Res}}^{B}_{A^{\tilde{\omega}}})(g\circ q_{c})$
$={\operatorname{Ho}}({\operatorname{Res}}^{B}_{A^{\tilde{\omega}}})(g)\circ{\operatorname{Ho}}({\operatorname{Res}}^{B}_{A^{\tilde{\omega}}})(q_{c}).$
Since $(-)\circ{\operatorname{Ho}}({\operatorname{Res}}^{B}_{A^{\tilde{\omega}}})(q_{c}):[{\mathfrak{m}}_{1},{\mathfrak{m}}_{2}]_{A^{\tilde{\omega}}}\to[Q_{c}{\mathfrak{m}}_{1},{\mathfrak{m}}_{2}]_{A^{\tilde{\omega}}}$ is an isomorphism, we deduce that ${\operatorname{Ho}}({\operatorname{Res}}^{B}_{A})(g)\cdot\tilde{\omega}={\operatorname{Ho}}({\operatorname{Res}}^{B}_{A^{\tilde{\omega}}})(g)$ as required. ∎

### v4

@@ 5.2. Supersingular characters in ${\operatorname{Ho}}({\mathcal{H}_{\operatorname{aff}}})$
The maps $\varepsilon_{\mathfrak{m}}:{\operatorname{Res}}^{\mathcal{H}}_{{\mathcal{H}}_{x}}({\mathfrak{m}})\otimes_{{\mathcal{H}}^{\dagger}_{x}}{\mathcal{H}}\to{\mathfrak{m}}$ and $\varepsilon_{{\mathfrak{m}}^{\prime}}:{\operatorname{Res}}^{\mathcal{H}}_{{\mathcal{H}}_{x}}({\mathfrak{m}}^{\prime})\otimes_{{\mathcal{H}}^{\dagger}_{x}}{\mathcal{H}}\to{\mathfrak{m}}^{\prime}$ from the Ollivier-Schneider resolution are trivial fibrations, cf. 5.17(ii), so that it again suffices to show that ${\operatorname{Res}}^{\mathcal{H}}_{{\mathcal{H}}^{\dagger}_{x}}({\mathfrak{m}})\cong{\operatorname{Res}}^{\mathcal{H}}_{{\mathcal{H}}^{\dagger}_{x}}({\mathfrak{m}}^{\prime})$ in ${\operatorname{Ho}}({\mathcal{H}}^{\dagger}_{x})$. Note that $\Omega_{x}=\langle\omega^{3}\rangle$. Similarly to the previous example we have decompositions ${\operatorname{Res}}^{\mathcal{H}}_{{\mathcal{H}}^{\dagger}_{x}}({\mathfrak{m}})\cong{\mathfrak{m}}_{0}\oplus{\mathfrak{m}}_{1}\oplus{\mathfrak{m}}_{2}$ and ${\operatorname{Res}}^{\mathcal{H}}_{{\mathcal{H}}^{\dagger}_{x}}({\mathfrak{m}}^{\prime})\cong{\mathfrak{m}}^{\prime}_{0}\oplus{\mathfrak{m}}^{\prime}_{1}\oplus{\mathfrak{m}}^{\prime}_{2}$. Here, all ${\mathfrak{m}}_{j}$ and ${\mathfrak{m}}^{\prime}_{j}$ are one-dimensional, and ${\mathcal{H}}_{x}$ acts on ${\mathfrak{m}}_{0}$, ${\mathfrak{m}}_{1}$ and ${\mathfrak{m}}_{2}$ via the characters $(-1,-1,0)_{\xi}$, $(0,-1,-1)_{\xi}$ and $(-1,0,-1)_{\xi}$ respectively, and on ${\mathfrak{m}}^{\prime}_{0}$, ${\mathfrak{m}}^{\prime}_{1}$ and ${\mathfrak{m}}^{\prime}_{2}$ via the characters $(0,-1,0)_{\xi}$, $(0,0,-1)_{\xi}$ and $(-1,0,0)_{\xi}$ respectively. Furthermore, $\hat{\omega}^{3}$ acts as $\lambda{\operatorname{id}}$ on both ${\mathfrak{m}}$ and ${\mathfrak{m}}^{\prime}$. Thus we see that ${\mathfrak{m}}_{0}\cong{\mathfrak{m}}^{\prime}_{0}$ and ${\mathfrak{m}}_{2}\cong{\mathfrak{m}}^{\prime}_{1}$, and 5.12(ii) together with [21, Lemma 3.4] imply that ${\mathfrak{m}}_{1}$ and ${\mathfrak{m}}^{\prime}_{2}$ have finite projective dimension over ${\mathcal{H}}^{\dagger}_{x}$. Putting everything together, this gives
@@ 5.3. The $\widetilde{\Omega}$-action on morphisms
We recall the action of subgroups of $\Omega$ on ${\operatorname{Hom}}$ spaces for various subalgebras of ${\mathcal{H}}$ as in [3, §4.2]. Let $A,B\subseteq{\mathcal{H}}$ be two Gorenstein $k$-subalgebras such that $k[T(\mathbb{F}_{q})]\subseteq A\subseteq B$ and $B$ is free over $A$, both as a left and as a right module. Furthermore, we also let $\widetilde{\Omega}_{B}:=\widetilde{\Omega}\cap B$ and denote by $\widetilde{\Omega}_{A}$ the normaliser of $A$ in $\widetilde{\Omega}_{B}$. One example to have in mind is $A={\mathcal{H}_{\operatorname{aff}}}$, $\widetilde{\Omega}_{A}=\widetilde{\Omega}_{B}=\widetilde{\Omega}$ and $B={\mathcal{H}}$, but we will see other situations where this applies as well. We will denote by $\Omega_{A}$ the image of $\widetilde{\Omega}_{A}$ under the canonical surjection $\widetilde{\Omega}\twoheadrightarrow\Omega$. We also denote by ${\mathcal{H}}_{A,\Omega_{A}}$ the $k$-subalgebra of ${\mathcal{H}}$ which is generated by $A$ and $\{T_{\tilde{\omega}}\}_{\tilde{\omega}\in\widetilde{\Omega}_{A}}$. To simplify notation, for any ${\mathfrak{m}}\in{\operatorname{Mod}}(B)$ we will denote ${\operatorname{Res}}^{B}_{A}({\mathfrak{m}})$ by ${\mathfrak{m}}|_{A}$.
@@
We fix $\tilde{\omega}\in\widetilde{\Omega}_{B}$ and let $A^{\tilde{\omega}}:=T_{\tilde{\omega}^{-1}}AT_{\tilde{\omega}}\subseteq B$. We have $A^{\tilde{\omega}}=A$ whenever $\tilde{\omega}\in\widetilde{\Omega}_{A}$, and in general the map $a\mapsto T_{\tilde{\omega}^{-1}}aT_{\tilde{\omega}}$ defines an isomorphism $A\cong A^{\tilde{\omega}}$ of $k$-algebras. Thus $A^{\tilde{\omega}}$ is also Gorenstein. This isomorphism induces an exact equivalence of categories $\mathcal{F}_{\tilde{\omega}}:{\operatorname{Mod}}(A)\xrightarrow{\simeq}{\operatorname{Mod}}(A^{\tilde{\omega}})$, where $\mathfrak{n}^{\tilde{\omega}}:=\mathcal{F}_{\tilde{\omega}}(\mathfrak{n})$ is defined to have the same underlying $k$-vector space as $\mathfrak{n}$ with $A^{\tilde{\omega}}$-action given by $x\cdot(T_{\tilde{\omega}^{-1}}aT_{\tilde{\omega}}):=xa$ for all $a\in A$ and all $x\in\mathfrak{n}^{\tilde{\omega}}$. Observe further that if ${\mathfrak{m}}\in{\operatorname{Mod}}(B)$ then the map $x\mapsto xT_{\tilde{\omega}}$ defines a natural $A^{\tilde{\omega}}$-linear isomorphism $\nu_{\mathfrak{m}}:({\mathfrak{m}}|_{A})^{\tilde{\omega}}\xrightarrow{\cong}{\mathfrak{m}}|_{A^{\tilde{\omega}}}$, i.e. there is a natural isomorphism $\nu:\mathcal{F}_{\tilde{\omega}}\circ{\operatorname{Res}}^{B}_{A}\to{\operatorname{Res}}^{B}_{A^{\tilde{\omega}}}$.
@@
Suppose now that ${\mathfrak{m}}_{1},{\mathfrak{m}}_{2}\in{\operatorname{Mod}}(B)$ and let $f\in{\operatorname{Hom}}_{A}({\mathfrak{m}}_{1}|_{A},{\mathfrak{m}}_{2}|_{A})$. By considering the image of $f$ under $\mathcal{F}_{\tilde{\omega}}$ and pulling back along the isomorphisms $\nu_{{\mathfrak{m}}_{i}}$ ($i=1,2$), we obtain an $A^{\tilde{\omega}}$-linear map $f\cdot\tilde{\omega}:{\mathfrak{m}}_{1}|_{A^{\tilde{\omega}}}\to{\mathfrak{m}}_{2}|_{A^{\tilde{\omega}}}$ defined by $(f\cdot\tilde{\omega})(x)=f(xT_{\tilde{\omega}^{-1}})T_{\tilde{\omega}}$ for $x\in{\mathfrak{m}}_{1}$. Thus we have constructed a map
@@
The above construction extends to morphisms in the homotopy category as follows. Since the functor $\mathcal{F}_{\tilde{\omega}}$ is exact and preserves free and projective modules, it preserves all weak equivalences by 2.8(i). Since $B|_{A}$ and $B|_{A^{\tilde{\omega}}}\cong\mathcal{F}_{\tilde{\omega}}(B|_{A})$ are free, we see in the same way that ${\operatorname{Res}}^{B}_{A}$ and ${\operatorname{Res}}^{B}_{A^{\tilde{\omega}}}$ also preserve all weak equivalences. Thus, these functors descend to homotopy categories and we get a diagram
(5.11) ${{\operatorname{Ho}}(B)}$${{\operatorname{Ho}}(A)}$${{\operatorname{Ho}}(A^{\tilde{\omega}})}$$\scriptstyle{{\operatorname{Ho}}({\operatorname{Res}}^{B}_{A})}$$\scriptstyle{{\operatorname{Ho}}({\operatorname{Res}}^{B}_{A^{\tilde{\omega}}})}$$\scriptstyle{\mathcal{F}_{\tilde{\omega}}}$
which commutes up to natural isomorphism (given by $\nu$). For ${\mathfrak{m}}_{1},{\mathfrak{m}}_{2}\in{\operatorname{Mod}}(B)$ and $f\in[{\mathfrak{m}}_{1},{\mathfrak{m}}_{2}]_{A}$, we now define $f\cdot\tilde{\omega}$ to be the pullback of $\mathcal{F}_{\tilde{\omega}}(f)$ along $\nu$. Again, this defines an action of $\Omega_{A}$ on $[{\mathfrak{m}}_{1},{\mathfrak{m}}_{2}]_{A}$. From (5.11), we immediately have the following:
@@
Suppose that ${\mathfrak{m}}_{1},{\mathfrak{m}}_{2}\in{\operatorname{Mod}}(B)$ and let $\tilde{\omega}\in\widetilde{\Omega}_{B}$. Then the diagram
(5.12) ${{[{\mathfrak{m}}_{1},{\mathfrak{m}}_{2}]_{B}}}$${{[{\mathfrak{m}}_{1},{\mathfrak{m}}_{2}]_{A}}}$${{[{\mathfrak{m}}_{1},{\mathfrak{m}}_{2}]_{A^{\tilde{\omega}}}}}$$\scriptstyle{{\operatorname{Ho}}({\operatorname{Res}}^{B}_{A})}$$\scriptstyle{{\operatorname{Ho}}({\operatorname{Res}}^{B}_{A^{\tilde{\omega}}})}$$\scriptstyle{(-)\cdot\tilde{\omega}}$
@@
Remark 5.23.
  1. (i)
    Assume that ${\mathfrak{m}}_{1}$ is Gorenstein projective in ${\operatorname{Mod}}(B)$ and let $\theta:\mathfrak{p}\twoheadrightarrow{\mathfrak{m}}_{2}$ be a surjection from a projective $B$-module. Then by 2.14, we have a commutative diagram
[…]
    where the top row is exact. By construction, the dashed map $[{\mathfrak{m}}_{1},{\mathfrak{m}}_{2}]_{B}\to[{\mathfrak{m}}_{1},{\mathfrak{m}}_{2}]_{A}^{\Omega_{A}}$ is given by ${\operatorname{Ho}}({\operatorname{Res}}^{B}_{A})$ and we see explicitly the last part of the Lemma here.
  2. (ii)
    Take $A={\mathcal{H}_{\operatorname{aff}}}$ and $B={\mathcal{H}}$ so that $\Omega_{A}=\Omega$, and assume that $\mathbb{G}$ is semisimple and that $p$ does not divide $\left|\Omega\right|$. Then the functor of taking $\Omega$-invariants is exact and the bottom row of (5.13) is exact. Moreover, in that case the two leftmost vertical arrows in (5.13) are isomorphisms and it then follows from the 5-lemma that $[{\mathfrak{m}}_{1},{\mathfrak{m}}_{2}]_{\mathcal{H}}\cong[{\mathfrak{m}}_{1},{\mathfrak{m}}_{2}]_{\mathcal{H}_{\operatorname{aff}}}^{\Omega}$ for all ${\mathfrak{m}}_{1},{\mathfrak{m}}_{2}\in{\operatorname{Mod}}({\mathcal{H}})$ with ${\mathfrak{m}}_{1}$ Gorenstein projective. In particular, the functor ${\operatorname{Ho}}({\operatorname{Res}}^{\mathcal{H}}_{\mathcal{H}_{\operatorname{aff}}})$ is faithful in that case. In general though, letting $\mathfrak{c}=\ker(\theta)$ (with $\theta$ as above) and recalling that $[{\mathfrak{m}}_{1},{\mathfrak{m}}_{2}]_{\mathcal{H}}\cong{\operatorname{Ext}}^{1}_{\mathcal{H}}({\mathfrak{m}}_{1},\mathfrak{c})$ and $[{\mathfrak{m}}_{1},{\mathfrak{m}}_{2}]_{\mathcal{H}_{\operatorname{aff}}}^{\Omega}\cong{\operatorname{Ext}}^{1}_{\mathcal{H}_{\operatorname{aff}}}({\mathfrak{m}}_{1},\mathfrak{c})^{\Omega}$ (cf. 2.16), the exact sequence in [3, §4.2, eq.(4.4)] shows that the kernel of the map $[{\mathfrak{m}}_{1},{\mathfrak{m}}_{2}]_{B}\to[{\mathfrak{m}}_{1},{\mathfrak{m}}_{2}]_{A}^{\Omega_{A}}$ is $H^{1}(\Omega,{\operatorname{Hom}}_{\mathcal{H}_{\operatorname{aff}}}({\mathfrak{m}}_{1},\mathfrak{c}))$ and hence may fail to be zero when $\Omega$ has non-trivial group cohomology. In general, ${\operatorname{Ho}}({\operatorname{Res}}^{\mathcal{H}}_{\mathcal{H}_{\operatorname{aff}}})$ may thus fail to be faithful.